Assume that A and B are matrices of the same size. Determine an expression for A if 2A − B = 5(A + 2B).


1
Expert's answer
2021-05-11T13:56:29-0400

"Solution: Given ~that, A ~and~ B~ are~ matrices ~of~ the ~same~ size. Addition ~and ~subtraction
\\of ~the ~ given ~ matrices ~are ~possible.
\\\therefore 2A − B = 5(A + 2B)
\\~~2A-B=5A+10B
\\~2A-5A=10B+B
\\~~~~~~~-3A=11B
\\~~~~~~~~ ~~\therefore A=-\frac{11}{3} B
\\Which ~is ~ the~ expression ~for~ A."